### v1

# Perfect Divisibility and Coloring of Some Bull-Free Graphs

Abstract
A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. A bull is a graph consisting of a triangle with two disjoint pendant edge, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is nonadjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set.
Chudnovsky and Sivaraman [J. Graph Theory 90 (2019) 54-60] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [The Electron. J. of Combin. 29 (2022) P3.19.] proved that every (fork, bull)-free graph is perfectly divisible. Chen and Xu [Discrete Appl. Math. 372 (2025) 298-307.] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. Let $H\in${{odd balloon}, $\{P_{8},C_{5}\}\}$. In this paper, we prove that every ($H$, bull)-free graph is perfectly divisible. We also prove that a ($P_{6}$, bull)-free graph is perfectly divisible if and only if it contains no Mycielski-Grötzsch graph as an induced subgraph. As corollaries, these graphs is $\binom{\omega+1}{2}-$colorable. Notice that every odd torch contains an odd hole, a $P_{5}$, and a fork. Therefore, our results generalize the results of these scholars. Moreover, we prove that every ($P_{6}$, bull)-free graph $G$ satisfies $\chi(G)\leq\omega(G)^{7}$.
Key words and phrases: bull-free graphs; perfectly divisibility; chromatic number.
AMS 2000 Subject Classifications: 05C15, 05C75
1 introduction
In this paper, all graphs are finite and simple. As usual, let $P_{k}$ and $C_{k}$ be a path and a cycle on $k$ vertices respectively. We say that a graph $G$ contains a graph $H$ if $H$ is isomorphic to an induced subgraph of $G$. A graph $G$ is $H$-free if it does not contain $H$. Analogously, for a family $\mathcal{H}$ of graphs, we say that $G$ is $\mathcal{H}$-free if $G$ induces no member of $\mathcal{H}$.
A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow\{1,2,...,k\}$ such that $f(u)\neq f(v)$ whenever $uv\in E$. We say that $G$ is k-colorable if $G$ admits a $k$-coloring. The chromatic number of $G$, denoted by $\chi(G)$, is the smallest positive integer $k$ such that $G$ is $k$-colorable. A clique (resp. stable set) of $G$ is a set of pairwise adjacent (resp. nonadjacent) vertices in $G$. The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$. For a given positive integer $k$, we use the notation $[k]$ to denote the set $\{1,\ldots,k\}$.
[…]
An induced cycle of length $k\geq 4$ is called a hole, and $k$ is the length of the hole. A hole is odd if $k$ is odd, and even otherwise. An antihole is the complement graph of a hole. A graph $G$ is said to be perfect if $\chi(H)=\omega(H)$ for every induced subgraph $H$ of $G$. The famous Strong Perfect Graph Theorem [7] was established by Chudnovsky et al. in 2006:
Theorem 1.1
[7] A graph $G$ is perfect if and only if $G$ is (odd hole, odd antihole)-free.
Let $F$ denote the Mycielski-Grötzsch graph $G_{4}$. A bull is a graph consisting of a triangle with two disjoint pendant edges, an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is nonadjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and a E is a graph obtained from an induced $P_{5}=v_{1}v_{2}v_{3}v_{4}v_{5}$ by adding an new vertex exactly adjacent to $v_{3}$ in the $P_{5}$. (See Figure 1)
Figure 1: Illustration of some graphs.
A graph is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. This concept was proposed by Hoáng in [12]. Chudnovsky and Sivaraman [9] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [17] proved that every (fork, bull)-free graph is perfectly divisible. Chen and Xu [5] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. In this paper, we prove the following theorems.
Theorem 1.2
Let $G$ be a (oddtorch,bull)(odd~torch,bull)-free graph. Then $G$ is perfectly divisible.
Theorem 1.3
Let $G$ be a $(P_{8},C_{5},bull)$-free graph. Then $G$ is perfectly divisible.
It is easy to verify that every odd torch contains an odd hole, a $P_{5}$, a fork, and the graph $E$. Therefore, all the class of (odd hole, bull)-free graphs, ($P_{5}$, bull)-free graphs, (fork, bull)-free, and ($E$, bull)-free graphs are subclasses of the class of (odd balloon, bull)-free graphs. Theorem 1.2 generalizes the results of Chudnovsky and Sivaraman [9] and Karthick et al. [17]; and Theorem 1.3 generalizes the result of Chen and Xu [5].
Recall that we use $F$ to denote the Mycielski-Grötzsch graph. It is certain that $F$ is ($P_{6}$, bull)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$.). In fact, the graph $G$ obtained from $F$ by adding a $K_{n}$ in which each vertex is adjacent to all the vertices in $F$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible. Since every ($P_{5}$, bull)-free graph is perfectly divisible, a natural problem is that under what conditions is ($P_{6}$, bull)-free graph perfectly divisible. In this paper, we prove the following theorem.
Theorem 1.4
Let $G$ be a $(P_{6},\text{bull})$-free graph. Then $G$ is perfectly divisible if and only if $G$ is $F$-free.
[…]
By a simple induction on $\omega(G)$, we have that $\chi(G)\leq\binom{\omega(G)+1}{2}$ for each perfectly divisible graph $G$. According to Theorems 1.2, 1.3 and 1.4, we can directly derive the following corollary.
Corollary 1.1
Let $G$ be a $(H,\text{bull})$-free graph, where $H\in\{\{odd~torch\},\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
Corollary 1.2
Let $G$ be a (odd torch, $C_{3}$)-free graph. Then $\chi(G)\leq 3$.
We also consider the binding function of ($P_{6}$, bull)-free graphs, and prove the following theorem.
Theorem 1.5
Let $G$ be a $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{7}$.
We will prove Theorem 1.2 in Section 3, prove Theorems 1.4 and 1.3 in Section 4, and prove Theorem 1.5 in Section 5.
2 Notations and Preliminary Results
For $X\subseteq V(G)$, we use $G[X]$ to denote the subgraph of $G$ induced by $X$. Let $v\in V(G)$, $X\subseteq V(G)$. We use $N_{G}(v)$ to denote the set of vertices adjacent to $v$. Let $d_{G}(v)=|N_{G}(v)|$, $M_{G}(v)=V(G)\setminus(N_{G}(v)\cup\{v\})$, $N_{G}(X)=\{u\in V(G)\setminus X\;|\;u$ has a neighbor in X}X\}, and $M_{G}(X)=V(G)\setminus(X\cup N_{G}(X))$. If it does not cause any confusion, we usually omit the subscript $G$ and simply write $N(v)$, $d(v)$, $M(v)$, $N(X)$ and $M(X)$.
For a subset $A$ of $V(G)$ and a vertex $b\in V(G)\setminus A$, we say that $b$ is complete to $A$ if $b$ is adjacent to every vertex of $A$, and that $b$ is anticomplete to $A$ if $b$ is not adjacent to any vertex of $A$. For two disjoint subsets $A$ and $B$ of $V(G)$, AA is complete to $B$ if every vertex of $A$ is complete to $B$, and $A$ is anticomplete to $B$ if every vertex of $A$ is anticomplete to $B$.
For $A,B\subseteq V(G)$, let $N_{A}(B)=N(B)\cap A$ and $M_{A}(B)=A\setminus(N_{A}(B)\cup B)$. For $u,v\in V(G)$, we simply write $u\sim v$ if $uv\in E(G)$, and write $u\not\sim v$ if $uv\not\in E(G)$.
3 Proof of Theorem 1.2
In this section, we will prove Theorem 1.2. Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$. We say that $S$ is a homogeneous set of $G$ if for any vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$. We say that a graph $G$ is locally perfect if for every vertex $v\in V(G)$, $G[N(v)]$ is perfect. The following useful lemmas is important to our proof.
Lemma 3.1
[14] Every minimal nonperfectly divisible graph has no homogeneous set.
Lemma 3.2
[8] If $G$ is a bull-free graph, then either $G$ has a homogeneous set or for every $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
Lemma 3.3
Let $G$ be a connected locally perfect bull-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
Proof. Suppose to its contrary that $H$ is an odd antihole of $G[M(v)]$ with $V(H)=\{v_{1},v_{2},...,v_{k}\}$, where $k\geq 7$ and $\overline{H}=v_{1}v_{2}\cdots v_{k}v_{1}$. Let $v^{\prime}\in N(v)$. We first prove that
if $y\in N(V(H))$, then $|N(y)\cap V(H)|\geq 2$. (1)
Suppose not, and we may by symmetry assume that $N(y)\cap V(H)=\{v_{1}\}$. But then $\{v_{1},v_{3},v_{5},y,v_{2}\}$ induces a bull, a contradiction. This proves (1).
We next prove that
if $x\in M(V(H))$ and $y\in N(V(H))$ such that $x\sim y$, then $N(y)\cap V(H)$ is a stable set. (2)
Suppose not, and without loss of generality, we may suppose that $v_{1},v_{n}\in N(y)\cap V(H)$ with $v_{1}v_{n}\in E(G)$, where $3\leq n\leq k-2$. We next will show that $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. If not, then let $2\leq n^{\prime}\leq n-1$ be the minimum integer such that $y\not\sim v_{n^{\prime}}$. If $n=3$, then $n^{\prime}=2$. To avoid an induced bull on $\{v_{1},v_{3},y,x,v_{4}\}$, we have that $y\sim v_{4}$; but then $\{v_{1},v_{4},y,x,v_{2}\}$ induces a bull, a contradiction. Hence, $n\geq 4$, and thus $v_{n}\sim v_{2}$ and $v_{n-1}\sim v_{1}$.
We can deduce that $n^{\prime}\neq 2$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{2}\}$; and $n^{\prime}\neq n-1$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{n-1}\}$. We have that $3\leq n^{\prime}\leq n-2$, and so $v_{n^{\prime}}\sim v_{n}$ and $v_{1}\sim v_{n^{\prime}-1}$ by the minimality of $n^{\prime}$. But then $\{v_{n^{\prime}-1},v_{n},y,x,v_{n^{\prime}},\}$ induces a bull, a contradiction. Therefore, $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. Similarly, with the same arguments, we can deduce that $y$ is complete to $\{v_{1},v_{k},v_{k-1},\cdots,v_{n}\}$, and this implies that $y$ is complete to $V(H)$, which contradicts to that $G$ is locally perfect. This proves (2).
Since $G$ is connected and $V(H)\subseteq M(v)$, there exist two adjacent vertices $x\in M(V(H))$ and $y\in N(V(H))$. By (1) and (2), we may by symmetry assume that $N(y)\cap V(H)=\{v_{1},v_{2}\}$. But then, $\{v_{k-3},v_{k-1},v_{k},v_{1},y\}$ induces a bull, a contradiction. This completes the proof of Lemma 3.3. $\square$
Lemma 3.4
Let $G$ be a minimal nonperfect divisible bull-free graph. Then $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
Proof. It is certain that $G$ is connected. Suppose to its contrary that there exists a vertex $v\in V(G)$ such that $G[M(v)]$ is perfect. But then, $G[\{v\}\cup M(v)]$ is perfect and $G[N(v)]$ has clique number less than $\omega(G)$, which contradicts to that $G$ is minimal nonperfectly divisible. Therefore, for every $v\in V(G)$, $G[M(v)]$ is imperfect. Moreover, since $G$ is minimal nonperfectly divisible, $G$ has no homogeneous set by Lemma 3.1. And so, for every $v\in V(G)$, $G[N(v)]$ is perfect by Lemma 3.2. This proves Lemma 3.4. $\square$
By Lemmas 3.3 and 3.4, we have that
Lemma 3.5
Let $G$ be a minimal nonperfectly divisible bull-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
we now prove Theorem 1.2.
Proof of Theorems 1.2: Suppose the Theorem 1.2 is not true. Let $G$ be a minimal nonperfectly divisible (odd torch, bull)-free graph. By Lemma 3.4, we have that $G$ is connected, and
for every vertex $v\in V(G)$, $G[N(v)]$ is perfect and $G[M(v)]$ is imperfect. (3)
Let $v\in V(G)$. By (3), Lemma 3.5 and Theorem 1.1, $G[M(v)]$ contains an odd hole, say $C=v_{1}v_{2}\cdots v_{k}v_{1}$, where $k\geq 5$ is an odd integer. Since $G$ is connected, there exist two adjacent vertices $x\in M(V(C))$ and $y\in N(V(C))$. Since $G$ is odd torch-free, we have that $N(y)\cap V(C)$ is not a stable set. Hence, we may by symmetry assume that $\{v_{1},v_{2}\}\subseteq N(y)\cap V(C)$.
By (3), we have that $y$ is not complete to $V(C)$ as $G[N(y)]$ is perfect. Let $3\leq n\leq k$ be the minimum integer such that $y\not\sim v_{n}$. But then $\{v_{n-2},v_{n-1},y,x,v_{n}\}$ induces a bull, a contradiction. This completes the proof of Theorem 1.2. $\square$
4 Proof of Theorems 1.3 and 1.4
In this section, we will prove Theorems 1.4 and 1.3. The following useful lemmas is important to our proof.
Lemma 4.1
Let $G$ be a connected locally perfect $(P_{6},bull)$-free graph with $\omega(G)\geq 3$. Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 5-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{1}$ be a 5-hole in $G[M(v)]$. From now on, the subscript is modulo 5 in the proof of the Lemma. For $1\leq i\leq 5$, let
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
$W_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2},v_{i+3}\}\};$
Let $X=\bigcup_{i=1}^{5}X_{i}$, $Y=\bigcup_{i=1}^{5}Y_{i}$, $Z=\bigcup_{i=1}^{5}Z_{i}$ and $W=\bigcup_{i=1}^{5}W_{i}$. We next prove the following claim.
Claim 4.1
$N(V(C))=X\cup Y\cup Z\cup W$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z\cup W$. Let $u\in N(V(C))$. Without loss of generality, suppose $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done. So, suppose $u$ has a neighbor in $\{v_{2},v_{3},v_{4},v_{5}\}$.
If $u$ is anticomplete to $\{v_{2},v_{5}\}$, then $x$ is not complete to $\{v_{3},v_{4}\}$ as otherwise $\{u,v_{3},v_{4},v_{2},v_{5}\}$ induces a bull. Hence, $u$ has exactly one neighbor in $\{v_{3},v_{4}\}$. Now, $u\in Y_{1}$ if $u\sim v_{3}$ and $u\in Y_{4}$ if $u\sim v_{4}$. Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume $u\sim v_{2}$.
To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$. By symmetry, we may assume $u\sim v_{3}$. If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$, and we are done. So, we may suppose $u$ has a neighbor in $\{v_{4},v_{5}\}$. Since $G$ is locally perfect, $u$ is not complete to $V(C)$. Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$. This proves Claim 4.1. $\square$
Consequently, we prove that
for every $u\in N(v)$, $u\in Y$. (4)
Since $G$ is connected, we claim that $u$ has a neighbor in $V(C)$. On the contrary, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}\subseteq M(V(C))$ and $x_{3}\in N(V(C))$. By symmetry, suppose $x_{3}\sim v_{1}$. By Claim 4.1, without loss of generality, we may assume that $x_{3}\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $x_{3}\in W_{1}$, then $\{x_{3},v_{3},v_{4},x_{2},v_{5}\}$ induces a bull; if $x_{3}\in X_{1}$, then there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{1}v_{2}v_{3}$; otherwise, there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{3}v_{4}v_{5}$. All are contradictions. This implies that $N(u)\cap V(C)\neq\mbox{{\rm\O }}$.
By Claim 4.1, without loss of generality, we assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $u\in X_{1}$, then there is an induced $P_{6}=vuv_{1}v_{2}v_{3}v_{4}$. If $u\in Z_{1}$, then $\{v,u,v_{2},v_{3},v_{4}\}$ induces a bull. If $u\in W_{1}$, then $\{u,v_{3},v_{4},v,v_{5}\}$ induces a bull. So, $u\in Y_{1}$. This proves (4).
We next prove that
if $x,y\in N(v)$ such that $x\sim y$, then $N_{C}(x)=N_{C}(y)$. (5)
Assume for contradiction that $N_{C}(x)\neq N_{C}(y)$. Without loss of generality, suppose $N_{C}(x)=\{v_{1},v_{3}\}$ by (4). If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull. If $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v_{3},x,y,v_{2},v_{5}\}$ induces a bull. By symmetry, in all other cases an induced bull also occurs. Hence, $N_{C}(x)=N_{C}(y)$. This proves (5).
Since $\omega(G)\geq 3$ and $v$ belongs to a maximum clique of $G$, there exist two adjacent vertices in $N(v)$. Let $S$ be a component of $N(v)$ which contains an edge. It is certain that $1<|V(S)|<|V(G)|$. We next prove that $V(S)$ is a homogeneous set of $G$. By (4) and (5), we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
Obviously, we have that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$. Suppose $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a nonneighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
To avoid an induced bull on $\{x,y,z,v_{1},v_{2}\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{5}\}$. So $z\in N(V(C))$. By Claim 4.1, $z$ has a neighbor in $\{v_{1},v_{3}\}$. By symmetry, we may assume that $z\sim v_{1}$. But then $\{v_{1},x,z,v,v_{4}\}$ induces a bull, a contradiction. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.1. $\square$
The proof of following lemma is similar with the proof of Lemma 4.1.
Lemma 4.2
Let $G$ be a connected locally perfect $(P_{8},C_{5},bull)$-free graph with $\omega(G)\geq 3$. Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 7-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. From now on, the subscript is modulo 7 in the proof of the Lemma. For $1\leq i\leq 7$, let
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
Let $X=\bigcup_{i=1}^{7}X_{i}$, $Y=\bigcup_{i=1}^{7}Y_{i}$ and $Z=\bigcup_{i=1}^{7}Z_{i}$. We next prove the following claim.
Claim 4.2
$N(V(C))=X\cup Y\cup Z$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z$. Let $u\in N(V(C))$. Without loss of generality, suppose $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done. So, suppose $u$ has a neighbor in $\{v_{2},v_{3},\cdots,v_{7}\}$.
[…]
Without loss of generality, $u\sim v_{2}$. To avoid an induced bull on $\{v_{1},v_{2},u,v_{3},v_{7}\}$, we may by symmetry assume that $u\sim v_{3}$. If $u\sim v_{4}$, then $u\sim v_{5}$ to forbid an induced bull on $\{v_{3},v_{4},u,v_{1},v_{5}\}$; similarly, $u\sim v_{7}$. Since $G$ is locally perfect, $u$ is not complete to $V(C)$, and so $u\not\sim v_{6}$; but then $\{v_{4},v_{5},u,v_{2},v_{6}\}$ induces a bull, a contradiction. So $u\not\sim v_{4}$, and similarly, $u\not\sim v_{7}$.
[…]
Claim 4.3
For every $u\in N(v)$, $u\in Y$. And for $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
Proof. Let $u\in N(v)$. If $N_{C}(u)=\mbox{{\rm\O }}$, since $G$ is connected, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}$ is anticomplete to $V(C)$ and $x_{3}$ has a neighbor in $V(C)$. We may by symmetry assume that $x_{3}\in X_{3}\cup Y_{1}\cup Z_{1}$ by Claim 4.2; but then there exists an induced $P_{8}=x_{!}x_{2}x_{3}v_{3}v_{4}v_{5}v_{6}v_{7}$. So $u\in N(V(C))$. By symmetry and Claim 4.2, we may assume that $u\in X_{3}\cup Y_{1}\cup Z_{1}$. If $u\in X_{1}$, then there is an induced $P_{8}=vuv_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$. If $u\in Z_{1}$, then $\{v_{2},v_{3},u,v,v_{4}\}$ induces a bull. So $u\in Y_{1}\subseteq Y$.
Let $x,y\in N(v)$ such that $x\sim y$. By symmetry, we may suppose that $N_{C}(x)=\{v_{1},v_{3}\}$ and $N_{C}(y)\in\{\{v_{1},v_{3}\},\{v_{2},v_{4}\},\{v_{3},v_{5}\},\{v_{4},v_{6}\}\}$. If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull; if $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v,x,y,v_{1},v_{5}\}$ induces a bull; if $N_{C}(y)=\{v_{4},v_{6}\}$, then $\{v,x,y,v_{1},v_{6}\}$ induces a bull. So, $N_{C}(y)=N_{C}(x)=\{v_{1},v_{3}\}$. This proves Claim 4.3. $\square$
Recall that $v$ belongs to a maximum clique of $G$ and $\omega(G)\geq 3$. Let $S$ be a component of $N(v)$ which contains an edge. It is certain that $1<|V(S)|<|V(G)|$. We next prove that $V(S)$ is a homogeneous set of $G$. By Claim 4.3, we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup(V(C)\setminus\{v_{1},v_{3}\})$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a nonneighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$. To avoid an induced bull on $\{x,y,v_{1},v_{2},z\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{7}\}$. So $z\in N(V(C))$, but $z\notin X\cup Y\cup Z$, which contradicts with Claim 4.2. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.2. $\square$
We need the following lemmas.
Lemma 4.3
[19] Let $G$ be a $(P_{6},C_{3})$-free graphs. Then $\chi(G)\leq 4$, and $G$ contains Mycielski-Grötzsch graph if $\chi(G)=4$.
Lemma 4.4
[10] Let $G$ be a $(P_{8},C_{5},C_{3})$-free graphs. Then $\chi(G)\leq 3$.
Now, we can prove Theorems 1.3 and 1.4.
Proof of Theorems 1.3 and 1.4: It suffices to prove that every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly divisible. Suppose not, and let $G$ be a minimal nonperfectly divisible ($H$, bull)-free graph, where $H\in\{P_{6},\{P_{8},C_{5}\}\}$. By Lemmas 3.4, we have that $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect. By Lemmas 4.3 and 4.4, we have that $\omega(G)\geq 3$ because $G$ is nonperfectly divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Lemma 3.5, $G[M(v)]$ is odd antihole-free except $C_{5}$. By Theorem 1.1, we have that $G[M(v)]$ contains an odd hole. Therefore, by Lemmas 4.1 and 4.2, $G$ has a homogeneous set, which contradicts with Lemma 3.1. This completes the proof of Theorems 1.3 and 1.4. $\square$
5 Proof of Theorem 1.5
In this section, we will prove Theorem 1.5. Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$. Let $\cal C$} be a class of graphs. We denote by ${\cal C}_{s}$ the closure of $\cal C$} under taking substitution.
We say a bull-free graph $G$ is basic if neither $G$ nor $\overline{G}$ contains an odd hole with both a vertex which is complete to the odd hole and a vertex which is anticomplete to the odd hole. Chudnovsky and Safra [8] proved the following two lemmas.
Lemma 5.1
((1.4 of [8] )) Every bull-free graph can be obtained via substitution from basic bull-free graphs.
Lemma 5.2
((4.3 of [8] )) Let $G$ be a basic bull-free graph. Then for every vertex $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
Lemma 5.3
((6.1 of [2] )) Let $\cal C$} be a class of graphs. If $\cal C$} is hereditary and Polynomially $\chi$-bounded with $\chi\leq\omega^{k}$, then ${\cal C}_{s}$ is polynomially $\chi$-bounded with $\chi\leq\omega^{2k+3}$.
Lemma 5.4
[18] The class of perfect graphs is closed under taking substitution.
Lemma 5.5
Let $G$ be a locally perfect $(P_{6},bull)$-free graph. Then there exists a partition $(X,Y)$ of $V(G)$ such that $G[X]$ and $G[Y]$ are perfect.
Proof. We prove it by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. If $\omega(G)\leq 2$, by Lemma 4.3, then $\chi(G)\leq 4$ and we are done. Therefore, by inductive hypothesis, we may assume that
$G$ is connected and $\omega(G)\geq 3$. (6)
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ is perfect, then $(N(v),M(v)\cup\{v\})$ is a partition of $V(G)$ which satisfies the Lemma since $G$ is locally perfect. Therefore, we may assume that $G[M(v)]$ is imperfect. By Lemma 3.3, $G[M(v)]$ contains no odd antihole of number of vertices at least 7. So, by Theorem 1.1, $G[M(v)]$ contains a 5-hole as $G$ is $P_{6}$-free. By Lemma 4.1, we have that $G$ has a homogeneous set, say $S$. Recall that $1<|S|<|V(G)|$.
Since $G$ is connected, we have that there exists a vertex which is complete to $S$, and so $G[S]$ is perfect because $G$ is locally perfect. Since $S\neq\mbox{{\rm\O }}$, we may assume that $x\in S$. Let $G^{\prime}=G[V(G)\setminus(S\setminus\{x\})]$. It is easy to verify that $G^{\prime}$ is a locally perfect $(P_{6}$, bull)-free graph and $|V(G^{\prime})|<|V(G)|$. Therefore, by inductive hypothesis, there exists a partition $(X^{\prime},Y^{\prime})$ of $V(G^{\prime})$ such that $G[X^{\prime}]$ and $G[Y^{\prime}]$ is perfect. Since $x\in V(G^{\prime})$, without loss of generality, we may assume that $x\in X^{\prime}$. Let $X=X^{\prime}\cup S$ and $Y=Y^{\prime}$. Clearly, $(X,Y)$ is a partition of $V(G)$. Since $S$ is a homogeneous set of $G$, we have that $S$ is a homogeneous set of $G[X]$, and thus $G[X]$ is a graph obtained from $G[X^{\prime}]$ by substituting the vertex $x$ by the graph $G[S]$. By Lemma 5.4, $G[X]$ is perfect. Therefore, the partition $(X,Y)$ satisfies the Lemma. This completes the proof of Lemma 5.5. $\square$
Lemma 5.6
Let $G$ be a basic $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{2}$.
Proof. We prove it by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. We may assume that $\omega(G)\geq 2$ and $G$ is connected by inductive hypothesis.
If $G$ is locally perfectly, then $\chi(G)\leq 2\omega(G)\leq\omega(G)^{2}$ by Lemma 5.5, and so we are done. Therefore, we may assume that $G$ is not locally perfect. Now, by Lemma 5.1, there exists a vertex $v$ such that $G[M(v)]$ is perfect. By inductive hypothesis, we have that $\chi(G[N(v)])\leq(\omega(G)-1)^{2}$. Therefore, we have that
$\chi(G)\leq\chi(G[N(v)])+\chi(G[\{v\}\cup M(v)])\leq(\omega(G)-1)^{2}+\omega(G)\leq\omega(G)^{2}.$
This completes the proof of Lemma 5.6. $\square$
Now, we can prove Theorem 1.5.
Proof of Theorem 1.5: Let $\cal G$} be the class of ($P_{6}$, bull)-free graphs, and $\cal C$} be the class of basic ($P_{6}$, bull)-free graphs. It is easy to verify that $\cal C$} is hereditary. By Lemmas 5.3 and 5.6, it suffices to show that ${\cal G}\subseteq{\cal C}_{s}$.
Let $G\in{\cal G}$. Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a basic bull-free graph $G_{1}$ by substituting a vertex by a basic bull-free graph $G_{2}$. Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraph of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free; this implies that $G_{1},G_{2}\in{\cal C}$. Therefore, $G\in{\cal C}_{s}$, and so ${\cal G}\subseteq{\cal C}_{s}$.
This completes the proof of Theorem 1.5. $\square$
Remark
[…]
More generally, let ${\cal G}_{1}$ be a hereditary class of graphs in which every trianlge-free graph is $k$-colorable, let ${\cal G}_{2}$ be the class of bull-free graphs, and let ${\cal G}={\cal G}_{1}\cap{\cal G}_{2}$. In [6], Chudnovsky et al. proved that $\cal G$} is polynomially $\chi$-bounded. Here, what are we more concerned about is that is it true that every graph in $\cal G$} is perfectly divisible if and only if $k=3$.
Declarations
  • Research of the first author was supported by Postgraduate Research and Practice Innovation Program of Jiangsu Province KYCX25_1926. And research of the second author was supported by the Scientific Research Foundation of Nanjing Institute of Technology, China (No. YKJ202448).
  • Conflict of interest The authors declare no conflict of interest.
  • Data availibility statement This manuscript has no associated data.