### v3

# Perfect Divisibility and Coloring of Some Bull-Free Graphs

Abstract
A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. A bull is a graph consisting of a triangle with two disjoint pendant edges, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set.
Chudnovsky and Sivaraman [J. Graph Theory 90 (2019) 54-60] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [The Electron. J. of Combin. 29 (2022) P3.19.] proved that every (fork, bull)-free graph is perfectly divisible. Chen and Xu [Discrete Appl. Math. 372 (2025) 298-307.] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. Let $H\in${{odd torch}, $\{P_{8},C_{5}\}\}$. In this paper, we prove that every ($H$, bull)-free graph is perfectly divisible. We also prove that a ($P_{6}$, bull)-free graph is perfectly divisible if and only if it contains no Mycielski-Grötzsch graph as an induced subgraph. As corollaries, these graphs are $\binom{\omega+1}{2}-$colorable. Notice that every odd torch contains an odd hole, an induced $P_{5}$, and an induced fork. Therefore, our results generalize their findings. Moreover, we prove that every ($P_{6}$, bull)-free graph $G$ satisfies $\chi(G)\leq\omega(G)^{7}$.
Key words and phrases: bull-free graphs; perfectly divisibility; chromatic number.
AMS 2000 Subject Classifications: 05C15, 05C75
1 introduction
In this paper, all graphs are finite and simple. We follow [1] for undefined notations and terminology. As usual, let $P_{k}$ and $C_{k}$ be a path and a cycle on $k$ vertices respectively. We say that a graph $G$ contains a graph $H$ if $H$ is isomorphic to an induced subgraph of $G$. A graph $G$ is $H$-free if it does not contain $H$. Analogously, for a family $\mathcal{H}$ of graphs, we say that $G$ is $\mathcal{H}$-free if $G$ induces no member of $\mathcal{H}$.
For a given positive integer $k$, we use the notation $[k]$ to denote the set $\{1,\ldots,k\}$. A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow[k]$ such that $f(u)\neq f(v)$ whenever $uv\in E$. We say that $G$ is k-colorable if $G$ admits a $k$-coloring. The chromatic number of $G$, denoted by $\chi(G)$, is the smallest positive integer $k$ such that $G$ is $k$-colorable. A clique (resp. stable set) of $G$ is a set of pairwise adjacent (resp. non-adjacent) vertices in $G$. The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$.
[…]
An induced cycle of length $k\geq 4$ is called a hole, and $k$ is the length of the hole. A hole is odd if $k$ is odd, and even otherwise. An antihole is the complement graph of a hole. A graph $G$ is said to be perfect if $\chi(H)=\omega(H)$ for every induced subgraph $H$ of $G$. The famous Strong Perfect Graph Theorem [7] was established by Chudnovsky et al. in 2006:
Theorem 1.1
[7] A graph $G$ is perfect if and only if $G$ is ((odd hole, odd antihole))-free.
Let $F$ denote the Mycielski-Grötzsch graph $G_{4}$. A bull is a graph consisting of a triangle with two disjoint pendant edges, an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and a E is a graph obtained from an induced $P_{5}=v_{1}v_{2}v_{3}v_{4}v_{5}$ by adding an new vertex exactly adjacent to $v_{3}$ in the $P_{5}$. (See Figure 1)
Figure 1: Illustration of some graphs.
A graph is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. This concept was proposed by Hoáng in [12]. Chudnovsky and Sivaraman [9] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [15] proved that every (fork, bull)-free graph is perfectly divisible, and this class of graphs was generalized by Wu and Xu [19]. Chen and Xu [5] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. In this paper, we prove the following theorems.
Theorem 1.2
Let $G$ be a (oddtorch,bull)(odd~torch,bull)-free graph. Then $G$ is perfectly divisible.
Theorem 1.3
Let $G$ be a $(P_{8},C_{5},bull)$-free graph. Then $G$ is perfectly divisible.
It is easy to verify that every odd torch contains an odd hole, an induced $P_{5}$, an induced fork, and the graph $E$. Therefore, all the class of (odd hole, bull)-free graphs, ($P_{5}$, bull)-free graphs, (fork, bull)-free, and ($E$, bull)-free graphs are subclasses of the class of (odd torch, bull)-free graphs. Theorem 1.2 generalizes the results of Chudnovsky and Sivaraman [9] and Karthick et al. [15]; and Theorem 1.3 generalizes the result of Chen and Xu [5].
Recall that we use $F$ to denote the Mycielski-Grötzsch graph. It is certain that $F$ is ($P_{6}$, bull)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$.). In fact, the graph $G$ obtained from $F$ by adding a $K_{n}$ in which each vertex is adjacent to all the vertices in $F$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible. Since every ($P_{5}$, bull)-free graph is perfectly divisible, a natural problem is that under what conditions is ($P_{6}$, bull)-free graph perfectly divisible ? In this paper, we prove the following theorem.
Theorem 1.4
Let $G$ be a $(P_{6},\text{bull})$-free graph. Then $G$ is perfectly divisible if and only if $G$ is $F$-free.
[…]
By a simple induction on $\omega(G)$, we have that $\chi(G)\leq\binom{\omega(G)+1}{2}$ for each perfectly divisible graph $G$. According to Theorems 1.2, 1.3 and 1.4, we can directly derive the following corollary.
Corollary 1.1
Let $G$ be an $(H,\text{bull})$-free graph, where $H\in\{\{odd~torch\},\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
Corollary 1.2
Let $G$ be an ((odd torch, $C_{3}$))-free graph. Then $\chi(G)\leq 3$.
We also consider the binding function of ($P_{6}$, bull)-free graphs, and prove the following theorem.
Theorem 1.5
Let $G$ be a $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{7}$.
Let $G$ be a graph and let $w$ be a weight function on $V(G)$. We use $\omega_{w}(G)$ to denote the maximum weight of a clique in $G$. A graph $G$ is perfectly weight divisible [9, 12] if for every positive integer weight function $w$ on $V(G)$, and for every induced subgraph $H$ of $G$, there is a partition of $V(H)$ into $P$, $W$ such that $H[P]$ is perfect and $\omega_{w}(H[W])<\omega_{w}(H)$. It is certain that if a graph $G$ is perfectly weight divisible, then $G$ is perfectly divisible. If $G$ is not perfectly weight divisible but each of its proper induced subgraph is, then we say $G$ is minimally non-perfectly weight divisible (MNWD for short).
As mentioned above, the structure and coloring problem of bull-free graphs has garnered significant attention, with new methods and research content [4, 6, 11, 13, 18] being continuously proposed. This paper utilizes the concept of perfectly weight divisibility to further analyze the structure of bull-free graphs and bull-free MNWD graphs, thereby deriving our main results. We will prove Theorem 1.2 in Section 3, prove Theorems 1.4 and 1.3 in Section 4, and prove Theorem 1.5 in Section 5.
2 Notations and Preliminary Results
For $X\subseteq V(G)$, we use $G[X]$ to denote the subgraph of $G$ induced by $X$. Let $v\in V(G)$, $X\subseteq V(G)$. We use $N_{G}(v)$ to denote the set of vertices adjacent to $v$. Let $d_{G}(v)=|N_{G}(v)|$, $M_{G}(v)=V(G)\setminus(N_{G}(v)\cup\{v\})$, $N_{G}(X)=\{u\in V(G)\setminus X\;|\;u$ has a neighbor in X}X\}, and $M_{G}(X)=V(G)\setminus(X\cup N_{G}(X))$. If it does not cause any confusion, we usually omit the subscript $G$ and simply write $N(v)$, $d(v)$, $M(v)$, $N(X)$ and $M(X)$.
For a subset $A$ of $V(G)$ and a vertex $b\in V(G)\setminus A$, we say that $b$ is complete to $A$ if $b$ is adjacent to every vertex of $A$, and that $b$ is anticomplete to $A$ if $b$ is not adjacent to any vertex of $A$. For two disjoint subsets $A$ and $B$ of $V(G)$, AA is complete to $B$ if every vertex of $A$ is complete to $B$, and $A$ is anticomplete to $B$ if every vertex of $A$ is anticomplete to $B$.
For $A,B\subseteq V(G)$, let $N_{A}(B)=N(B)\cap A$ and $M_{A}(B)=A\setminus(N_{A}(B)\cup B)$. For $u,v\in V(G)$, we simply write $u\sim v$ if $uv\in E(G)$, and write $u\not\sim v$ if $uv\not\in E(G)$.
3 Proof of Theorem 1.2
Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$. We say that $S$ is a homogeneous set of $G$ if every vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$. We say that a graph $G$ is locally perfect if for every vertex $v\in V(G)$, $G[N(v)]$ is perfect.
In this section, we will prove Theorem 1.2. The following useful lemmas is important to our proof.
Lemma 3.1
[9] Every MNWD graph has no homogeneous set.
Lemma 3.2
[8] If $G$ is a bull-free graph, then either $G$ has a homogeneous set or for every $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
Lemma 3.3
Let $G$ be a connected locally perfect bull-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
Proof. Suppose to its contrary that $H$ is an odd antihole of $G[M(v)]$ with $V(H)=\{v_{1},v_{2},...,v_{k}\}$, where $k\geq 7$ and $\overline{H}=v_{1}v_{2}\cdots v_{k}v_{1}$. Let $v^{\prime}\in N(v)$. We first prove that
if $y\in N(V(H))$, then $|N(y)\cap V(H)|\geq 2$. (1)
Suppose not, and we may by symmetry assume that $N(y)\cap V(H)=\{v_{1}\}$. But then $\{v_{1},v_{3},v_{5},y,v_{2}\}$ induces a bull, a contradiction. This proves (1).
We next prove that
if $x\in M(V(H))$ and $y\in N(V(H))$ such that $x\sim y$, then $N(y)\cap V(H)$ is a stable set. (2)
Suppose not, and without loss of generality, we may suppose that $v_{1},v_{n}\in N(y)\cap V(H)$ with $v_{1}v_{n}\in E(G)$, where $3\leq n\leq k-2$. We next show that $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. If not, then let $2\leq n^{\prime}\leq n-1$ be the minimum integer such that $y\not\sim v_{n^{\prime}}$. If $n=3$, then $n^{\prime}=2$. To avoid an induced bull on $\{v_{1},v_{3},y,x,v_{4}\}$, we have that $y\sim v_{4}$; but then $\{v_{1},v_{4},y,x,v_{2}\}$ induces a bull, a contradiction. Hence, $n\geq 4$, and thus $v_{n}\sim v_{2}$ and $v_{n-1}\sim v_{1}$.
We can deduce that $n^{\prime}\neq 2$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{2}\}$; and $n^{\prime}\neq n-1$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{n-1}\}$. We have that $3\leq n^{\prime}\leq n-2$, and so $v_{n^{\prime}}\sim v_{n}$ and $v_{1}\sim v_{n^{\prime}-1}$ by the minimality of $n^{\prime}$. But then $\{v_{n^{\prime}-1},v_{n},y,x,v_{n^{\prime}},\}$ induces a bull, a contradiction. Therefore, $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. Similarly, with the same arguments, we can deduce that $y$ is complete to $\{v_{1},v_{k},v_{k-1},\cdots,v_{n}\}$, and this implies that $y$ is complete to $V(H)$, which contradicts to that $G$ is locally perfect. This proves (2).
Since $G$ is connected and $V(H)\subseteq M(v)$, there exist two adjacent vertices $x\in M(V(H))$ and $y\in N(V(H))$. By (1) and (2), we may by symmetry assume that $N(y)\cap V(H)=\{v_{1},v_{2}\}$. But then, $\{v_{k-3},v_{k-1},v_{k},v_{1},y\}$ induces a bull, a contradiction. This completes the proof of Lemma 3.3. $\square$
Lemma 3.4
Let $G$ be a bull-free MNWD graph. Then $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
Proof. It is certain that $G$ is connected. Since $G$ is MNWD, we have that for every vertex $v\in V(G)$, $G[M(v)]$ cannot be perfect as otherwise, $(\{v\}\cup M(v),N(v))$ is a partition of $V(G)$ such that for all positive integer weight function on $V(G)$, the $G[\{v\}\cup M(v)]$ is a perfect graph and $G[N(v)]$ has a smaller maximum weight of a clique than $G$, which contradicts with the minimality of $G$. Therefore, for every $v\in V(G)$, $G[M(v)]$ is imperfect. Moreover, since $G$ is MNWD, $G$ has no homogeneous set by Lemma 3.1. Therefore, for every $v\in V(G)$, $G[N(v)]$ is perfect by Lemma 3.2. This proves Lemma 3.4. $\square$
By Lemmas 3.3 and 3.4, we have that
Lemma 3.5
Let $G$ be a bull-free MNWD graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
We now prove Theorem 1.2.
Proof of Theorems 1.2: It suffices to show that each (odd torch, bull)-free graph is perfectly weight divisible. Suppose it is not true, and let $G$ be a (odd torch, bull)-free MNWD graph. By Lemma 3.4, we have that $G$ is connected, and
for every vertex $v\in V(G)$, $G[N(v)]$ is perfect and $G[M(v)]$ is imperfect. (3)
Let $v\in V(G)$. By (3), Lemma 3.5 and Theorem 1.1, $G[M(v)]$ contains an odd hole, say $C=v_{1}v_{2}\cdots v_{k}v_{1}$, where $k\geq 5$ is an odd integer. Since $G$ is connected, there exist two adjacent vertices $x\in M(V(C))$ and $y\in N(V(C))$. Since $G$ is odd torch-free, we have that $N(y)\cap V(C)$ is not a stable set. Hence, we may by symmetry assume that $\{v_{1},v_{2}\}\subseteq N(y)\cap V(C)$.
By (3), we have that $y$ is not complete to $V(C)$ as $G[N(y)]$ is perfect. Let $3\leq n\leq k$ be the minimum integer such that $y\not\sim v_{n}$. But then $\{v_{n-2},v_{n-1},y,x,v_{n}\}$ induces a bull, a contradiction. Therefore, each (odd torch, bull)-free graph is perfectly weight divisible. This completes the proof of Theorem 1.2. $\square$
4 Proof of Theorems 1.3 and 1.4
In this section, we will prove Theorems 1.3 and 1.4. The following useful lemmas are important to our proof.
Lemma 4.1
Let $G$ be a graph with $\chi(G)\leq 3$. Then $G$ is perfectly weight divisible.
Proof. Let $w$ be a positive integer weight function on $V(G)$, and let $G^{\prime}$ be an induced subgraph of $G$. Since $\chi(G^{\prime})\leq 3$, there is a partition $(A,B)$ of $V(G^{\prime})$ such that $G^{\prime}[A]$ is bipartite and $B$ is a stable set of $G^{\prime}$. Let $X\subseteq B$ such that each vertex in $X$ is a isolated vertex of $G^{\prime}$. Then $(A\cup X,B\setminus X)$ is a partition of $G^{\prime}$ such that $G[A\cup X]$ is perfect, and $\omega_{w}(G[B\setminus X])<\omega_{w}(G^{\prime})$. This proves the lemma. $\square$
Lemma 4.2
Let $G$ be a connected locally perfect $(P_{6},bull)$-free graph with $\omega(G)\geq 3$. Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 5-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{1}$ be a 5-hole in $G[M(v)]$. From now on, the subscript is modulo 5 in the proof of the Lemma. For $1\leq i\leq 5$, let
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
$W_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2},v_{i+3}\}\};$
Let $X=\bigcup_{i=1}^{5}X_{i}$, $Y=\bigcup_{i=1}^{5}Y_{i}$, $Z=\bigcup_{i=1}^{5}Z_{i}$ and $W=\bigcup_{i=1}^{5}W_{i}$. We next prove the following claim.
Claim 4.1
$N(V(C))=X\cup Y\cup Z\cup W$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z\cup W$. Let $u\in N(V(C))$. Without loss of generality, suppose $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done. So, suppose $u$ has a neighbor in $\{v_{2},v_{3},v_{4},v_{5}\}$.
If $u$ is anticomplete to $\{v_{2},v_{5}\}$, then $x$ is not complete to $\{v_{3},v_{4}\}$ as otherwise $\{u,v_{3},v_{4},v_{2},v_{5}\}$ induces a bull. Hence, $u$ has exactly one neighbor in $\{v_{3},v_{4}\}$. Now, $u\in Y_{1}$ if $u\sim v_{3}$ and $u\in Y_{4}$ if $u\sim v_{4}$. Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume that $u\sim v_{2}$.
To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$. By symmetry, we may assume that $u\sim v_{3}$. If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$, and we are done. So, we may suppose that $u$ has a neighbor in $\{v_{4},v_{5}\}$. Since $G$ is locally perfect, we have that $u$ is not complete to $V(C)$. Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$. This proves Claim 4.1. $\square$
Consequently, we prove that
for every $u\in N(v)$, $u\in Y$. (4)
Since $G$ is connected, we claim that $u$ has a neighbor in $V(C)$. On the contrary, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}\subseteq M(V(C))$ and $x_{3}\in N(V(C))$. By symmetry, suppose $x_{3}\sim v_{1}$. By Claim 4.1, without loss of generality, we may assume that $x_{3}\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $x_{3}\in W_{1}$, then $\{x_{3},v_{3},v_{4},x_{2},v_{5}\}$ induces a bull; if $x_{3}\in X_{1}$, then there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{1}v_{2}v_{3}$; otherwise, there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{3}v_{4}v_{5}$. All are contradictions. This implies that $N(u)\cap V(C)\neq\mbox{{\rm\O }}$.
By Claim 4.1, without loss of generality, we assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $u\in X_{1}$, then there is an induced $P_{6}=vuv_{1}v_{2}v_{3}v_{4}$. If $u\in Z_{1}$, then $\{v,u,v_{2},v_{3},v_{4}\}$ induces a bull. If $u\in W_{1}$, then $\{u,v_{3},v_{4},v,v_{5}\}$ induces a bull. So, $u\in Y_{1}$. This proves (4).
We next prove that
if $x,y\in N(v)$ such that $x\sim y$, then $N_{C}(x)=N_{C}(y)$. (5)
Assume for contradiction that $N_{C}(x)\neq N_{C}(y)$. Without loss of generality, suppose $N_{C}(x)=\{v_{1},v_{3}\}$ by (4). If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull. If $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v_{3},x,y,v_{2},v_{5}\}$ induces a bull. By symmetry, in all other cases an induced bull also occurs. Hence, $N_{C}(x)=N_{C}(y)$. This proves (5).
Since $\omega(G)\geq 3$ and $v$ belongs to a maximum clique of $G$, there exist two adjacent vertices in $N(v)$. Let $S$ be a component of $N(v)$ which contains an edge. It is certain that $1<|V(S)|<|V(G)|$. We next prove that $V(S)$ is a homogeneous set of $G$. By (4) and (5), we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
Obviously, we have that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
To avoid an induced bull on $\{x,y,z,v_{1},v_{2}\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{5}\}$. So $z\in N(V(C))$. By Claim 4.1, $z$ has a neighbor in $\{v_{1},v_{3}\}$. By symmetry, we may assume that $z\sim v_{1}$. But then $\{v_{1},x,z,v,v_{4}\}$ induces a bull, a contradiction. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.2. $\square$
The proof of the following lemma is similar to that of Lemma 4.2.
Lemma 4.3
Let $G$ be a connected locally perfect $(P_{8},C_{5},bull)$-free graph with $\omega(G)\geq 3$. Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 7-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. For the remainder of the proof, all subscripts are taken modulo 7. For $1\leq i\leq 7$, let
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
Let $X=\bigcup_{i=1}^{7}X_{i}$, $Y=\bigcup_{i=1}^{7}Y_{i}$ and $Z=\bigcup_{i=1}^{7}Z_{i}$. We next prove the following claim.
Claim 4.2
$N(V(C))=X\cup Y\cup Z$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z$. Let $u\in N(V(C))$. Without loss of generality, suppose $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done. So, suppose $u$ has a neighbor in $\{v_{2},v_{3},\cdots,v_{7}\}$.
Suppose that $u$ is anticomplete to $\{v_{2},v_{7}\}$. Then $u$ has a neighbor in $\{v_{3},v_{4},v_{5},v_{6}\}$. Assume that $u$ is anticomplete to $\{v_{3},v_{6}\}$. If $u\sim v_{4}$, then there is a 5-hole $uv_{1}v_{2}v_{3}v_{4}u$, and we are done by Lemma 4.2. So, $u\not\sim v_{4}$; similarly, $u\not\sim v_{5}$, a contradiction. Therefore, $u$ is not anticomplete to $\{v_{3},v_{6}\}$ and we may by symmetry assume that $u\sim v_{3}$. If $u\sim v_{6}$, then $u$ has a neighbor in $\{v_{4},v_{5}\}$ to avoid a 5-hole $uv_{3}v_{4}v_{5}v_{6}u$; and we may by symmetry assume that $u\sim v_{4}$, but now $\{v_{3},v_{4},u,v_{2},v_{6}\}$ induces a bull. So, $u\not\sim v_{6}$. To avoid a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$, we have that $u\not\sim v_{5}$. Moreover, to avoid an induced bull on $\{v_{3},v_{4},u,v_{2},v_{5}\}$, we have that $u\not\sim v_{4}$. Now, $N(u)\cap V(C)=\{v_{1},v_{3}\}$, then $u\in Y_{1}$ and we are done. So, $u$ is not anticomplete to $\{v_{2},v_{7}\}$.
Without loss of generality, $u\sim v_{2}$. To avoid an induced bull on $\{v_{1},v_{2},u,v_{3},v_{7}\}$, we may by symmetry assume that $u\sim v_{3}$. If $u\sim v_{4}$, then $u\sim v_{5}$ to forbid an induced bull on $\{v_{3},v_{4},u,v_{1},v_{5}\}$; similarly, $u\sim v_{7}$. Since $G$ is locally perfect, $u$ is not complete to $V(C)$, and so $u\not\sim v_{6}$; but then $\{v_{4},v_{5},u,v_{2},v_{6}\}$ induces a bull, a contradiction. So $u\not\sim v_{4}$, and similarly, $u\not\sim v_{7}$.
If $u$ is complete to $\{v_{5},v_{6}\}$, then $\{v_{5},v_{6},u,v_{4},v_{7}\}$ induces a bull. So $u$ has at most one neighbor in $\{v_{5},v_{6}\}$. If $u\sim v_{5}$, then there is a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$. Hence, $u$ is anticomplete to $\{v_{5},v_{6}\}$ by symmetry. And thus $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$; this implies $u\in Z_{1}$. This proves Claim 4.2. $\square$
Claim 4.3
For every $u\in N(v)$, we have that $u\in Y$. And for $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
Proof. Let $u\in N(v)$. If $N_{C}(u)=\mbox{{\rm\O }}$, since $G$ is connected, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}$ is anticomplete to $V(C)$ and $x_{3}$ has a neighbor in $V(C)$. We may by symmetry assume that $x_{3}\in X_{3}\cup Y_{1}\cup Z_{1}$ by Claim 4.2; but then there exists an induced $P_{8}=x_{!}x_{2}x_{3}v_{3}v_{4}v_{5}v_{6}v_{7}$. So $u\in N(V(C))$. By symmetry and Claim 4.2, we may assume that $u\in X_{3}\cup Y_{1}\cup Z_{1}$. If $u\in X_{1}$, then there is an induced $P_{8}=vuv_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$. If $u\in Z_{1}$, then $\{v_{2},v_{3},u,v,v_{4}\}$ induces a bull. So $u\in Y_{1}\subseteq Y$.
Let $x,y\in N(v)$ such that $x\sim y$. By symmetry, we may suppose that $N_{C}(x)=\{v_{1},v_{3}\}$ and $N_{C}(y)\in\{\{v_{1},v_{3}\},\{v_{2},v_{4}\},\{v_{3},v_{5}\},\{v_{4},v_{6}\}\}$. If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull; if $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v,x,y,v_{1},v_{5}\}$ induces a bull; if $N_{C}(y)=\{v_{4},v_{6}\}$, then $\{v,x,y,v_{1},v_{6}\}$ induces a bull. So, $N_{C}(y)=N_{C}(x)=\{v_{1},v_{3}\}$. This proves Claim 4.3. $\square$
Recall that $v$ belongs to a maximum clique of $G$ and $\omega(G)\geq 3$. Let $S$ be a component of $N(v)$ which contains an edge. It is certain that $1<|V(S)|<|V(G)|$. We next prove that $V(S)$ is a homogeneous set of $G$. By Claim 4.3, we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup(V(C)\setminus\{v_{1},v_{3}\})$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$. To avoid an induced bull on $\{x,y,v_{1},v_{2},z\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{7}\}$. So $z\in N(V(C))$, but $z\notin X\cup Y\cup Z$, which contradicts with Claim 4.2. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.3. $\square$
We need the following lemmas.
Lemma 4.4
[17] Let $G$ be a $(P_{6},C_{3})$-free graphs. Then $\chi(G)\leq 4$, and $G$ contains Mycielski-Grötzsch graph if $\chi(G)=4$.
Lemma 4.5
[10] Let $G$ be a $(P_{8},C_{5},C_{3})$-free graphs. Then $\chi(G)\leq 3$.
Now, we can prove Theorems 1.3 and 1.4.
Proof of Theorems 1.3 and 1.4: Since the graph $F$ is non-perfectly divisible, it suffices to prove that every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly divisible. We will actually prove the following stronger conclusion: every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly weight divisible. Suppose not, and let $G$ be a ($H$, bull)-free MNWD graph, where $H\in\{P_{6},\{P_{8},C_{5}\}\}$. By Lemma 3.4, we have that $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect. By Lemmas 4.14.4 and 4.5, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Lemma 3.5, $G[M(v)]$ contains no odd antihole of number of vertices at least 7. By Theorem 1.1, we have that $G[M(v)]$ contains an odd hole. Therefore, by Lemmas 4.2 and 4.3, $G$ has a homogeneous set, which contradicts with Lemma 3.1. This completes the proof of Theorems 1.3 and 1.4. $\square$
5 Proof of Theorem 1.5
In this section, we will prove Theorem 1.5. A class of graphs is hereditary if it is closed under taking induced subgraphs. Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$. Let $\cal C$} be a class of graphs. We denote by ${\cal C}_{s}$ the closure of $\cal C$} under taking substitution.
We say a bull-free graph $G$ is basic if neither $G$ nor $\overline{G}$ contains an odd hole with both a vertex which is complete to the odd hole and a vertex which is anticomplete to the odd hole. Chudnovsky and Safra [8] proved the following two lemmas.
Lemma 5.1
((1.4 of [8] )) Every bull-free graph can be obtained via substitution from basic bull-free graphs.
Lemma 5.2
((4.3 of [8] )) Let $G$ be a basic bull-free graph. Then for every vertex $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
Lemma 5.3
((6.1 of [2] )) Let $\cal C$} be a class of graphs. If $\cal C$} is hereditary and polynomially $\chi$-bounded with $\chi\leq\omega^{k}$, then ${\cal C}_{s}$ is polynomially $\chi$-bounded with $\chi\leq\omega^{2k+3}$.
Lemma 5.4
[16] The class of perfect graphs is closed under taking substitution.
Lemma 5.5
Let $G$ be a locally perfect $(P_{6},bull)$-free graph. Then there exists a partition $(X,Y)$ of $V(G)$ such that $G[X]$ and $G[Y]$ are perfect.
Proof. We prove it by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. If $\omega(G)\leq 2$, by Lemma 4.4, then $\chi(G)\leq 4$ and we are done. Therefore, by inductive hypothesis, we may assume that
$G$ is connected and $\omega(G)\geq 3$. (6)
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ is perfect, then $(N(v),M(v)\cup\{v\})$ is a partition of $V(G)$ which satisfies the Lemma since $G$ is locally perfect. Therefore, we may assume that $G[M(v)]$ is imperfect. By Lemma 3.3, $G[M(v)]$ contains no odd antihole of number of vertices at least 7. So, by Theorem 1.1, $G[M(v)]$ contains a 5-hole as $G$ is $P_{6}$-free. By Lemma 4.2, we have that $G$ has a homogeneous set, say $S$. Recall that $1<|S|<|V(G)|$.
Since $G$ is connected, we have that there exists a vertex which is complete to $S$, and so $G[S]$ is perfect because $G$ is locally perfect. Since $S\neq\mbox{{\rm\O }}$, we may assume that $x\in S$. Let $G^{\prime}=G[V(G)\setminus(S\setminus\{x\})]$. It is easy to verify that $G^{\prime}$ is a locally perfect $(P_{6}$, bull)-free graph and $|V(G^{\prime})|<|V(G)|$. Therefore, by inductive hypothesis, there exists a partition $(X^{\prime},Y^{\prime})$ of $V(G^{\prime})$ such that $G[X^{\prime}]$ and $G[Y^{\prime}]$ is perfect. Since $x\in V(G^{\prime})$, without loss of generality, we may assume that $x\in X^{\prime}$. Let $X=X^{\prime}\cup S$ and $Y=Y^{\prime}$. Clearly, $(X,Y)$ is a partition of $V(G)$. Since $S$ is a homogeneous set of $G$, we have that $S$ is a homogeneous set of $G[X]$, and thus $G[X]$ is a graph obtained from $G[X^{\prime}]$ by substituting the vertex $x$ by the graph $G[S]$. By Lemma 5.4, $G[X]$ is perfect. Therefore, the partition $(X,Y)$ satisfies the lemma. This completes the proof of Lemma 5.5. $\square$
Lemma 5.6
Let $G$ be a basic $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{2}$.
Proof. We prove it by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. We may assume that $\omega(G)\geq 2$ and $G$ is connected by inductive hypothesis.
If $G$ is locally perfectly, then $\chi(G)\leq 2\omega(G)\leq\omega(G)^{2}$ by Lemma 5.5, and so we are done. Therefore, we may assume that $G$ is not locally perfect. Now, by Lemma 5.1, there exists a vertex $v$ such that $G[M(v)]$ is perfect. By inductive hypothesis, we have that $\chi(G[N(v)])\leq(\omega(G)-1)^{2}$. Therefore, we have that
$\chi(G)\leq\chi(G[N(v)])+\chi(G[\{v\}\cup M(v)])\leq(\omega(G)-1)^{2}+\omega(G)\leq\omega(G)^{2}.$
This completes the proof of Lemma 5.6. $\square$
Now, we can prove Theorem 1.5.
Proof of Theorem 1.5: Let $\cal G$} be the class of ($P_{6}$, bull)-free graphs, and $\cal C$} be the class of basic ($P_{6}$, bull)-free graphs. It is easy to verify that $\cal C$} is hereditary. By Lemmas 5.3 and 5.6, it suffices to show that ${\cal G}\subseteq{\cal C}_{s}$.
Let $G\in{\cal G}$. Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a basic bull-free graph $G_{1}$ by substituting a vertex by a basic bull-free graph $G_{2}$. Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraphs of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free, which implies that $G_{1},G_{2}\in{\cal C}$. Therefore, $G\in{\cal C}_{s}$, and so ${\cal G}\subseteq{\cal C}_{s}$.
This completes the proof of Theorem 1.5. $\square$
Remark
[…]
More generally, let ${\cal G}_{1}$ be a hereditary class of graphs in which every trianlge-free graph is $k$-colorable, let ${\cal G}_{2}$ be the class of bull-free graphs, and let ${\cal G}={\cal G}_{1}\cap{\cal G}_{2}$. In [6], Chudnovsky et al. proved that $\cal G$} is polynomially $\chi$-bounded. Here, what are we more concerned about is that is it true that every graph in $\cal G$} is perfectly divisible if and only if $k=3$ ?
Declarations
  • Ran Chen is supported by Postgraduate Research and Practice Innovation Program of Jiangsu Province KYCX25_1926. Di Wu is supported by the Scientific Research Foundation of Nanjing Institute of Technology, China (No. YKJ202448).
  • Conflict of interest The authors declare no conflict of interest.
  • Data availibility statement This manuscript has no associated data.