### v5

# On Perfect Divisibility of Bull-Free Graphs Without Long Paths

Abstract
A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. A bull is a graph consisting of a triangle with two disjoint pendant edges. Let $H$ be an odd hole or an odd antihole. An odd net is the graph obtained from $H$ by adding a new vertex adjacent to all vertices of $H$. An odd net+ is the graph obtained from an odd net by adding an additional new vertex that is non-adjacent to any vertex of $H$. Let $F$ denote the Grötzsch graph.
Chudnovsky and Sivaraman [J. Graph Theory 90 (2019) 54-60] proved (bull, $P_{5}$)-free graphs are perfectly divisible. Chen and Xu [Discrete Appl. Math. 372 (2025) 298-307] proved that (bull, $P_{7},C_{5}$)-free graphs are perfectly divisible. In this paper, we prove that (bull, $P_{8},C_{5}$)-free graphs are perfectly divisible. As known $F$ is non-perfectly divisible, and so the class of (bull, $P_{6}$)-free graphs is non-perfectly divisible. Let $G$ be a (bull, $P_{6}$)-free graph. We also show that (i) $G$ is perfectly divisible if and only if $G$ is $F$-free; (ii) if $G$ is odd net-free, then $G$ can be partitioned into two perfect graphs, and so $\chi(G)\leq 2\omega(G)$; (iii) if $G$ is odd net+-free, then $\chi(G)\leq\omega(G)^{2}$; (iv) $\chi(G)\leq\omega(G)^{7}$. These results generalize the findings of Chudnovsky and Sivaraman, and Chen and Xu.
Key words and phrases: bull-free graphs; perfect divisibility; chromatic number.
AMS 2020 Subject Classifications: 05C15, 05C75
1 introduction
In this paper, all graphs are finite and simple. We follow [1] for undefined notations and terminology. As usual, let $P_{k}$ and $C_{k}$ be a path and a cycle on $k$ vertices respectively. We say that a graph $G$ contains a graph $H$ if $H$ is isomorphic to an induced subgraph of $G$. A graph $G$ is $H$-free if it does not contain $H$. Analogously, for a family $\mathcal{H}$ of graphs, we say that $G$ is $\mathcal{H}$-free if $G$ induces no member of $\mathcal{H}$.
Let $H_{1}$ and $H_{2}$ be two vertex disjoint graphs. The join $H_{1}+H_{2}$ is the graph with $V(H_{1}+H_{2})=V(H_{1})\cup V(H_{2})$ and $E(H_{1}+H_{2})=E(H_{1})\cup E(H_{2})\cup\{xy\;|\;x\in V(H_{1}),y\in V(H_{2})$}\}.
For a given positive integer $k$, we use $[k]$ to denote the set $\{1,\ldots,k\}$. A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow[k]$ such that $f(u)\neq f(v)$ whenever $uv\in E$. We say that $G$ is k-colorable if $G$ admits a $k$-coloring. The chromatic number of $G$, denoted by $\chi(G)$, is the smallest positive integer $k$ such that $G$ is $k$-colorable. A clique (resp. stable set) of $G$ is a set of pairwise adjacent (resp. non-adjacent) vertices in $G$. The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$.
The concept of $\chi$-boundedness was introduced by Gyárfás [15] in 1975. A class of graphs is said to be hereditary if it is closed under isomorphism and induced subgraphs. For a hereditary class $\mathcal{F}$, if there exists a function $f$ such that $\chi(H)\leq f(\omega(H))$ for every $H\in\mathcal{F}$, then $\mathcal{F}$ is called $\chi$-bounded and $f$ a binding function.
An induced cycle of length $k\geq 4$ is called a hole, and $k$ is the length of the hole. A k-hole is a hole of length $k$. A hole is odd if $k$ is odd, and even otherwise. An antihole is the complement graph of a hole. A graph $G$ is said to be perfect if $\chi(H)=\omega(H)$ for every induced subgraph $H$ of $G$. The Strong Perfect Graph Theorem [8] was established by Chudnovsky et al. in 2006:
Theorem 1.1
[8] A graph $G$ is perfect if and only if $G$ is ((odd hole, odd antihole))-free.
Let $F$ denote the Grötzsch graph. A bull is a graph consisting of a triangle with two disjoint pendant edges. Let $H$ be an odd hole or an odd antihole. An odd net is the graph obtained from $H$ by adding a new vertex adjacent to all vertices of $H$. An odd net+ is the graph obtained from an odd net by adding an additional new vertex that is non-adjacent to any vertex of $H$. (see Figure 1).
Figure 1: Illustration of some forbidden graphs.
A graph $G$ is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. This concept was proposed by Hoàng in [13]. By a simple induction on $\omega(G)$, we have that every perfectly divisible graph $G$ satisfies $\chi(G)\leq\binom{\omega(G)+1}{2}$ [10].
Let $G$ be a graph and let $w$ be a weight function on $V(G)$. We use $\omega_{w}(G)$ to denote the maximum weight of a clique in $G$. A graph $G$ is perfectly weight divisible [10, 13] if for every positive integer weight function $w$ on $V(G)$, and for every induced subgraph $H$ of $G$, there is a partition of $V(H)$ into $P$ and $W$ such that $H[P]$ is perfect and $\omega_{w}(H[W])<\omega_{w}(H)$. If $G$ is not perfectly weight divisible but each of its proper induced subgraph is, then we say $G$ is minimally non-perfectly weight divisible (MNWD for short). It is certain that if a graph $G$ is perfectly weight divisible, then $G$ is perfectly divisible.
Chudnovsky and Sivaraman [10] proved that every (bull, odd hole)-free graph and every (bull, $P_{5}$)-free graph are perfectly divisible. Karthick et al. [16] proved that every (bull, fork)-free graph is perfectly divisible, where fork is the graph obtained from a $K_{1,3}$ by subdividing an edge once. Chen and Xu [5] proved that every (bull, $P_{7},C_{5}$)-free graph is perfectly divisible. It is certain that $F$ is (bull, $P_{6}$)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$). Actually, the graph $F+K_{n}$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible as $F$ is not perfectly divisible. Further properties of perfectly divisible graphs and bull-free graphs can be found in [6, 12, 14]. Since every (bull, $P_{5}$)-free graph is perfectly divisible, a natural question is to determine the conditions under which a (bull, $P_{6}$)-free graph is perfectly divisible.
In this paper, we generalize the results of Chudnovsky and Sivaraman [10], and Chen and Xu [5] by the following two theorems.
Theorem 1.2
Let $G$ be a $(bull,P_{8},C_{5})$-free graph. Then $G$ is perfectly weight divisible, and hence $G$ is perfectly divisible.
Theorem 1.3
Let $G$ be a $(bull,P_{6})$-free graph. Then $G$ is perfectly divisible if and only if $G$ is $F$-free.
Note that the class of $3K_{1}$-free graphs has no linear binding function [19]. In 2022, Brause et al.[3] gave a more general theorem.
Theorem 1.4
[…]
By Theorem 1.4, the classes of $(\text{bull},H)$-free graphs, where $H\in\{\{P_{8},C_{5}\},\{P_{6},F\}\}$, do not have a linear binding function. By Theorems 1.2 and 1.3, we can directly derive the following corollaries by a simple induction on $\omega(G)$.
Corollary 1.1
Let $G$ be an $(\text{bull},H)$-free graph, where $H\in\{\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
[…]
Theorem 1.5
Let $G$ be a ((bull, $P_{6}$, odd net))-free graph. Then $G$ can be partitioned into two perfect graphs, and hence $\chi(G)\leq 2\omega(G)$.
Theorem 1.6
[…]
Theorem 1.7
Let $G$ be a $(bull,P_{6})$-free graph. Then $\chi(G)\leq\omega(G)^{7}$.
We will prove Theorem 1.2 in Section 3, prove Theorem 1.3 in Section 4, and prove Theorems 1.5, 1.6 and 1.7 in Section 5.
2 Notations and Preliminary Results
For $X\subseteq V(G)$, we use $G[X]$ to denote the subgraph of $G$ induced by $X$. Let $v\in V(G)$, $X\subseteq V(G)$. We use $N_{G}(v)$ to denote the set of vertices adjacent to $v$. Let $d_{G}(v)=|N_{G}(v)|$, $M_{G}(v)=V(G)\setminus(N_{G}(v)\cup\{v\})$, $N_{G}(X)=\{u\in V(G)\setminus X\;|\;u$ has a neighbor in X}X\}, and $M_{G}(X)=V(G)\setminus(X\cup N_{G}(X))$. If it does not cause any confusion, we usually omit the subscript $G$ and simply write $N(v)$, $d(v)$, $M(v)$, $N(X)$ and $M(X)$.
For a subset $A$ of $V(G)$ and a vertex $u\in V(G)\setminus A$, we say that $u$ is complete to $A$ if $u$ is adjacent to every vertex of $A$, and that $u$ is anticomplete to $A$ if $u$ is not adjacent to any vertex of $A$. For two disjoint subsets $A$ and $B$ of $V(G)$, AA is complete to $B$ if every vertex of $A$ is complete to $B$, and $A$ is anticomplete to $B$ if every vertex of $A$ is anticomplete to $B$.
For $A,B\subseteq V(G)$, let $N_{A}(B)=N(B)\cap A$ and $M_{A}(B)=A\setminus(N_{A}(B)\cup B)$. For $u,v\in V(G)$, we simply write $u\sim v$ if $uv\in E(G)$, and write $u\not\sim v$ if $uv\not\in E(G)$.
Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$. We say that $S$ is a homogeneous set of $G$ if every vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$. The following useful lemmas are important to our proof.
Lemma 2.1
((Theorem 3.6 of [10])) Every MNWD graph has no homogeneous set.
Lemma 2.2
((4.3 of [9])) If $G$ is a bull-free graph, then either $G$ has a homogeneous set or for every $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
Lemma 2.3
((Lemma 6 of [4])) Let $G$ be a bull-free MNWD graph. Then $G$ is a connected odd net-free graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect and contains no odd antihole of number of vertices at least $7$.
Lemma 2.4
((Theorem 2 of [11])) Let $G$ be a $(P_{8},C_{5},C_{3})$-free graphs. Then $\chi(G)\leq 3$.
Lemma 2.5
((Theorem 6 of [18])) Let $G$ be a $(P_{6},C_{3})$-free graphs. Then $\chi(G)\leq 4$, and $G$ contains $F$ if $\chi(G)=4$.
Lemma 2.6
Let $G$ be a graph with $\chi(G)\leq 3$. Then $G$ is perfectly weight divisible.
Proof. Let $w$ be a positive integer weight function on $V(G)$, and let $G^{\prime}$ be an induced subgraph of $G$. Since $\chi(G^{\prime})\leq 3$, there is a partition $(A,B)$ of $V(G^{\prime})$ such that $G^{\prime}[A]$ is bipartite and $B$ is a stable set of $G^{\prime}$. Let $X\subseteq B$ such that each vertex in $X$ is an isolated vertex of $G^{\prime}$. Then $(A\cup X,B\setminus X)$ is a partition of $G^{\prime}$ such that $G[A\cup X]$ is perfect, and $\omega_{w}(G[B\setminus X])<\omega_{w}(G^{\prime})$. This proves the lemma. $\square$
3 Proof of Theorem 1.2
In this section, we prove Theorem 1.2 and begin with the following lemma.
Lemma 3.1
Let $G$ be a connected $(bull,P_{8},C_{5},odd~net)$-free graph with $\omega(G)\geq 3$, and let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a $7$-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. For the remainder of the proof, all subscripts are taken modulo 7. Let $i\in[7]$, define
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
Let $X=\bigcup_{i=1}^{7}X_{i}$, $Y=\bigcup_{i=1}^{7}Y_{i}$ and $Z=\bigcup_{i=1}^{7}Z_{i}$. We next prove the following claim.
Claim 3.1
$N(V(C))=X\cup Y\cup Z$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z$. Let $u\in N(V(C))$. Without loss of generality, suppose that $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$ and we are done. So we may assume that $u$ has a neighbor in $\{v_{2},v_{3},\cdots,v_{7}\}$.
Suppose that $u$ is anticomplete to $\{v_{2},v_{7}\}$. Then $u$ has a neighbor in $\{v_{3},v_{4},v_{5},v_{6}\}$. Assume that $u$ is anticomplete to $\{v_{3},v_{6}\}$. If $u\sim v_{4}$, then there is a 5-hole $uv_{1}v_{2}v_{3}v_{4}u$, a contradiction. So, $u\not\sim v_{4}$; similarly, $u\not\sim v_{5}$. Therefore, $u$ is not anticomplete to $\{v_{3},v_{6}\}$ and we may by symmetry assume that $u\sim v_{3}$. If $u\sim v_{6}$, then $u$ has a neighbor in $\{v_{4},v_{5}\}$ to avoid a 5-hole $uv_{3}v_{4}v_{5}v_{6}u$; and we may by symmetry assume that $u\sim v_{4}$, but now $\{v_{4},u,v_{3},v_{2},v_{6}\}$ induces a bull. So, $u\not\sim v_{6}$. To avoid a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$, we have that $u\not\sim v_{5}$. Moreover, to avoid an induced bull on $\{u,v_{3},v_{4},v_{2},v_{5}\}$, we have that $u\not\sim v_{4}$. Now, $N(u)\cap V(C)=\{v_{1},v_{3}\}$, and then $u\in Y_{1}$. So, we may assume that $u$ is not anticomplete to $\{v_{2},v_{7}\}$.
Without loss of generality, $u\sim v_{2}$. To avoid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{7}\}$, we may assume that $u\sim v_{3}$. If $u\sim v_{4}$, then $u\sim v_{5}$ to forbid an induced bull on $\{v_{3},u,v_{4},v_{1},v_{5}\}$; similarly, $u\sim v_{7}$. Since $G$ is odd net-free, $u$ is not complete to $V(C)$, and so $u\not\sim v_{6}$; but then $\{v_{4},v_{5},u,v_{2},v_{6}\}$ induces a bull, a contradiction. So $u\not\sim v_{4}$, and similarly, $u\not\sim v_{7}$.
If $u$ is complete to $\{v_{5},v_{6}\}$, then $\{u,v_{5},v_{6},v_{4},v_{7}\}$ induces a bull. So $u$ has at most one neighbor in $\{v_{5},v_{6}\}$. If $u\sim v_{5}$, then there exists a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$. So, $u\not\sim v_{5}$. Similarly, $u\not\sim v_{6}$. Then $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$; this implies $u\in Z_{1}$. This proves Claim 3.1. $\square$
Claim 3.2
If there exists $u\in N(v)$ such that $u\notin N(V(C))$, then $G$ has a homogeneous set.
Proof. Let $u\in N(v)$ such that $u\notin N(V(C))$. Then $u\in M(V(C))$. Since $G$ is connected, it follows that there exists an induced $P=x_{1}x_{2}\cdots x_{h}$ connecting $uv$ and $C$, where $h\geq 3$, such that $x_{1}\in\{u,v\}$ and $x_{h}\in V(C)$. Without loss of generality, $x_{1}=u$ and $x_{h}=v_{1}$. Since $x_{h-1}\in N(V(C))$, we may by symmetry assume that $x_{h-1}\in X_{1}\cup Y_{1}\cup Z_{1}$ by Claim 3.1. If $x_{h-1}\in X_{1}$, then there exists an induced $P_{8}=x_{h-2}x_{h-1}v_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$. If $x_{h-1}\in Z_{1}$, then $\{v_{2},v_{1},x_{h-1},v_{7},x_{h-2}\}$ induces a bull. Therefore $x_{h-1}\in Y_{1}$.
Given that $h=3$ as otherwise there exists an induced $P_{8}=v_{7}v_{6}v_{5}v_{4}v_{3}x_{h-1}x_{h-2}x_{h-3}$. So $x_{2}\sim v_{1}$ and $x_{2}\sim v_{3}$ as $x_{h-1}\in Y_{1}$. Consequently, to avoid an induced $P_{8}=v_{7}v_{6}v_{5}v_{4}v_{3}x_{2}x_{1}v$, we have that $v\sim x_{2}$. Let $S$ be the component of $G[N(x_{2})\cap M(V(C))]$ which contains the edge $uv$. We claim that $V(S)$ is a homogeneous set of $G$.
Suppose that there exists a vertex $t\in V(G)\setminus V(S)$ such that $t$ is neither complete nor anticomplete to $V(S)$. So there exist two vertices $t_{1},t_{2}\in V(S)$ such that $tt_{1}t_{2}$ is an induced $P_{3}$. Clearly $t\notin V(C)$ as $t_{1}\in M(V(C))$. We next show that
$t\not\sim x_{2}.$ (1)
Suppose to the contrary that $t\sim x_{2}$. By the choice of $S$, we have that $t\in N(V(C))$ as $t\notin V(S)$. Then $t$ has at least two neighbors in $V(C)$ because $t$ has a neighbor in $M(V(C))$ and $G$ is $P_{8}$-free. Hence $t\in Y\cup Z$ by Claim 3.1.
Suppose that $t\not\sim v_{1}$. Moreover, $t$ is anticomplete to $V(C)\setminus\{v_{2},v_{3},v_{7}\}$ as otherwise $\{t,t_{1},x_{2},v_{1},t^{\prime}\}$ induces a bull, where $t^{\prime}\in V(C)\setminus\{v_{1},v_{2},v_{3},v_{7}\}$ is a neighbor of $t$, a contradiction. If $t\sim v_{3}$, then $\{t,x_{2},v_{3},t_{2},v_{2}\}$ induces a bull whenever $t\not\sim v_{2}$, and $\{t,v_{2},v_{3},v_{1},v_{4}\}$ induces a bull whenever $t\sim v_{2}$. Both are contradictions. So $t\not\sim v_{3}$ and $t$ is anticomplete to $V(C)\setminus\{v_{2},v_{7}\}$. By Claim 3.1, $t\in Y_{7}$ as $t$ has at least two neighbors in $V(C)$. But then there exists an induced $P_{8}=v_{6}v_{5}v_{4}v_{3}v_{2}tt_{1}t_{2}$. So $t\sim v_{1}$, and by symmetry $t\sim v_{3}$.
Now $t\in Y_{1}\cup Z_{1}$ by Claim 3.1. But then $\{t,x_{2},v_{3},v_{4},t_{2}\}$ induces a bull. This proves (1).
By (1), we have that $t\sim v_{3}$ to avoid a bull on $\{t_{2},t_{1},x_{2},t,v_{3}\}$. Similarly, $t\sim v_{1}$. By Claim 3.1, $t\in Y_{1}\cup Z_{1}$. But then there exists an induced $P_{8}=v_{7}v_{6}v_{5}v_{4}v_{3}tt_{1}t_{2}$, a contradiction. Therefore $V(S)$ is a homogeneous set of $G$. This proves Claim 3.2. $\square$
By Claim 3.2, we may assume that
for every $u\in N(v)$, $u\in N(V(C))$. (2)
Claim 3.3
For every $u\in N(v)$, $u\in Y$. And for any $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
Proof. Let $u\in N(v)$. Recall that $G$ is connected. If $N_{C}(u)=\mbox{{\rm\O }}$, then there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}$ is anticomplete to $V(C)$ and $x_{3}$ has a neighbor in $V(C)$. We may by symmetry assume that $x_{3}\in X_{1}\cup Y_{1}\cup Z_{1}$ by Claim 3.1; but then there exists an induced $P_{8}=x_{1}x_{2}x_{3}v_{1}v_{7}v_{6}v_{5}v_{4}$. So $u\in N(V(C))$. Without loss of generality, we may assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}$. If $u\in X_{1}$, then there is an induced $P_{8}=vuv_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$. If $u\in Z_{1}$, then $\{v_{2},v_{3},u,v,v_{4}\}$ induces a bull. So $u\in Y_{1}\subseteq Y$.
Let $x,y\in N(v)$ such that $x\sim y$. By symmetry, we may suppose that $N_{C}(x)=\{v_{1},v_{3}\}$ and $N_{C}(y)\in\{\{v_{1},v_{3}\},\{v_{2},v_{4}\},\{v_{3},v_{5}\},\{v_{4},v_{6}\}\}$. If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull; if $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v,x,y,v_{1},v_{5}\}$ induces a bull; if $N_{C}(y)=\{v_{4},v_{6}\}$, then $\{v,x,y,v_{1},v_{6}\}$ induces a bull. So, $N_{C}(y)=N_{C}(x)=\{v_{1},v_{3}\}$. This proves Claim 3.3. $\square$
Recall that $v$ belongs to a maximum clique of $G$ and $\omega(G)\geq 3$. Let $S$ be a component of $N(v)$ which contains an edge. It is certain that $1<|V(S)|<|V(G)|$. We next prove that $V(S)$ is a homogeneous set of $G$. By Claim 3.3, we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup(V(C)\setminus\{v_{1},v_{3}\})$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exists an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$. To avoid an induced bull on $\{x,y,v_{1},v_{2},z\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{7}\}$. So $z\in N(V(C))$, but $z\notin X\cup Y\cup Z$, which contradicts with Claim 3.1. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 3.1. $\square$
We now proceed to prove Theorem 1.2.
Proof of Theorem 1.2: On the contrary, let $G$ be a (bull, $P_{8},C_{5}$)-free MNWD graph. By Lemma 2.3, we have that $G$ is a connected odd net-free graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect and contains no odd antihole of number of vertices at least 7. By Lemmas 2.4 and 2.6, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Theorem 1.1, we have that $G[M(v)]$ contains a 7-hole as $G$ is ($P_{8},C_{5}$)-free. Therefore, by Lemma 3.1, $G$ has a homogeneous set, which contradicts Lemma 2.1. This completes the proof of Theorem 1.2. $\square$
4 Proof of Theorem 1.3
In this section, we will prove Theorem 1.3. The following useful lemma is important to our proof, and its proof is similar to the proof of Lemma 3.1.
Lemma 4.1
Let $G$ be a connected $(bull,P_{6},odd~net)$-free graph with $\omega(G)\geq 3$, and let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 5-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{1}$ be a 5-hole in $G[M(v)]$. From now on, the subscript is modulo 5 in the proof of the lemma. For $i\in[5]$, let
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
$W_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2},v_{i+3}\}\};$
Let $X=\bigcup_{i=1}^{5}X_{i}$, $Y=\bigcup_{i=1}^{5}Y_{i}$, $Z=\bigcup_{i=1}^{5}Z_{i}$ and $W=\bigcup_{i=1}^{5}W_{i}$. We next prove the following claim.
Claim 4.1
$N(V(C))=X\cup Y\cup Z\cup W$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z\cup W$. Let $u\in N(V(C))$. Without loss of generality, suppose that $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$ and we are done. So, suppose that $u$ has a neighbor in $\{v_{2},v_{3},v_{4},v_{5}\}$.
If $u$ is anticomplete to $\{v_{2},v_{5}\}$, then $x$ is not complete to $\{v_{3},v_{4}\}$ as otherwise $\{u,v_{3},v_{4},v_{2},v_{5}\}$ induces a bull. Hence, $u$ has exactly one neighbor in $\{v_{3},v_{4}\}$. Now, $u\in Y_{1}$ if $u\sim v_{3}$ and $u\in Y_{4}$ if $u\sim v_{4}$. Therefore, we may assume that $u$ has a neighbor in $\{v_{2},v_{5}\}$, and by symmetry assume that $u\sim v_{2}$.
To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$. By symmetry, we may assume that $u\sim v_{3}$. If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$ and we are done. So, we may suppose that $u$ has a neighbor in $\{v_{4},v_{5}\}$. Since $G$ is odd net-free, we have that $u$ is not complete to $V(C)$. Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$. This proves Claim 4.1. $\square$
We next prove that
Claim 4.2
If there exists $u\in N(v)$ such that $u\notin N(V(C))$, then $G$ has a homogeneous set.
Proof. Let $u\in N(v)$ such that $u\notin N(V(C))$. Then $u\in M(V(C))$. Since $G$ is connected, it follows that there exists an induced $P=x_{1}x_{2}\cdots x_{h}$ connecting $uv$ and $C$, where $h\geq 3$, such that $x_{1}\in\{u,v\}$ and $x_{h}\in V(C)$. Without loss of generality, $x_{1}=u$ and $x_{h}=v_{1}$. Since $x_{h-1}\in N(V(C))$, we may by symmetry assume that $x_{h-1}\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$ by Claim 4.1. If $x_{h-1}\in X_{1}$, then there exists an induced $P_{6}=x_{h-2}x_{h-1}v_{1}v_{2}v_{3}v_{4}$. If $x_{h-1}\in Z_{1}\cup W_{1}$, then $\{v_{2},v_{1},x_{h-1},v_{5},x_{h-2}\}$ induces a bull. Therefore $x_{h-1}\in Y_{1}$.
Given that $h=3$ as otherwise there exists an induced $P_{6}=v_{5}v_{4}v_{3}x_{h-1}x_{h-2}x_{h-3}$. So $x_{2}\sim v_{1}$ and $x_{2}\sim v_{3}$ as $x_{h-1}\in Y_{1}$. Consequently, to avoid an induced $P_{6}=v_{5}v_{4}v_{3}x_{2}x_{1}v$, we have that $v\sim x_{2}$. Let $S$ be the component of $G[N(x_{2})\cap M(V(C))]$ which contains the edge $uv$. We claim that $V(S)$ is a homogeneous set of $G$.
Suppose to the contrary that there exists a vertex $t\in V(G)\setminus V(S)$ such that $t$ is neither complete nor anticomplete to $V(S)$. So there exist two vertices $t_{1},t_{2}\in V(S)$ such that $tt_{1}t_{2}$ is an induced $P_{3}$. Clearly $t\notin V(C)$ as $t_{1}\in M(V(C))$. We next show that
$t\not\sim x_{2}.$ (3)
Suppose to the contrary that $t\sim x_{2}$. By the choice of $S$, we have that $t\in N(V(C))$ as $t\notin V(S)$. It is certain that $t$ has at least two neighbors in $V(C)$ because $t$ has a neighbor in $M(V(C))$ and $G$ is $P_{6}$-free.
Suppose that $t\not\sim v_{1}$. Then $t\not\sim v_{4}$ as otherwise $\{t,t_{1},x_{2},v_{1},v_{4}\}$ induces a bull, a contradiction. Consequently, $t\not\sim v_{3}$ as otherwise $\{t,x_{2},v_{3},t_{2},v_{2}\}$ induces a bull whenever $t\not\sim v_{2}$, and $\{t,v_{2},v_{3},v_{1},v_{4}\}$ induces a bull whenever $t\sim v_{2}$. Both are contradictions. So $t$ is anticomplete to $\{v_{1},v_{3},v_{4}\}$. By Claim 4.1, $t\in Y_{5}$. But now there is an induced $P_{6}=v_{4}v_{3}v_{2}tt_{1}t_{2}$. Therefore $t\sim v_{1}$, and by symmetry $t\sim v_{3}$.
Now, $t\sim v_{2}$ as otherwise $\{t,v_{3},x_{2},v_{2},t_{2}\}$ induces a bull. By Claim 4.1, we may by symmetry assume that $t\in Z_{1}\cup W_{1}$. But then $\{v_{2},v_{1},t,v_{5},t_{1}\}$ induces a bull. This proves (3).
By (3), we have that $t\sim v_{3}$ to avoid a bull on $\{t_{1},t_{2},x_{2},t,v_{3}\}$. Similarly, $t\sim v_{1}$. By Claim 4.1, $t\in Y_{1}\cup W_{3}$. But then $v_{5}v_{4}v_{3}tt_{1}t_{2}$ is an induced $P_{6}$ if $t\in Y_{1}$, and $\{v_{5},v_{1},t,v_{2},t_{1}\}$ induces a bull if $t\in W_{3}$. Both are contradictions. Therefore $V(S)$ is a homogeneous set of $G$. This proves Claim 4.2. $\square$
By Claim 4.2, we may assume that for every $u\in N(v)$,
$u\in N(V(C)).$
Consequently, we prove that
for every $u\in N(v)$, $u\in Y$. (4)
By Claim 4.1, without loss of generality, we assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $u\in X_{1}$, then there is an induced $P_{6}=vuv_{1}v_{2}v_{3}v_{4}$. If $u\in Z_{1}$, then $\{v_{2},u,v_{3},v_{4},v\}$ induces a bull. If $u\in W_{1}$, then $\{v_{3},v_{4},u,v,v_{5}\}$ induces a bull. So, $u\in Y_{1}$. This proves (4).
We next prove that
if $x,y\in N(v)$ such that $x\sim y$, then $N_{C}(x)=N_{C}(y)$. (5)
Assume for contradiction that $N_{C}(x)\neq N_{C}(y)$. Without loss of generality, suppose that $N_{C}(x)=\{v_{1},v_{3}\}$ by (4). If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull. If $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v_{3},x,y,v_{2},v_{5}\}$ induces a bull. By symmetry, in all other cases, an induced bull also occurs. Hence, $N_{C}(x)=N_{C}(y)$. This proves (5).
Since $\omega(G)\geq 3$ and $v$ belongs to a maximum clique of $G$, there exist two adjacent vertices in $N(v)$. Let $S$ be a component of $N(v)$ which contains an edge. It is certain that $1<|V(S)|<|V(G)|$. We next prove that $V(S)$ is a homogeneous set of $G$. By (4) and (5), we may by symmetry assume that $N(V(S))\cap V(C)=\{v_{1},v_{3}\}$ as $S$ is connected.
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, it follows that there exists an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
To avoid an induced bull on $\{y,x,v_{1},z,v_{2}\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{5}\}$. So $z\in N(V(C))$. By Claim 4.1, $z$ has a neighbor in $\{v_{1},v_{3}\}$. By symmetry, we may assume that $z\sim v_{1}$. But then $\{v_{1},x,z,v,v_{4}\}$ induces a bull, a contradiction. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.1. $\square$
Now we begin to prove Theorem 1.3.
Proof of Theorem 1.3: Since the graph $F$ is non-perfectly divisible, it suffices to prove that every (bull, $P_{6}$, $F$)-free graph are perfectly weight divisible. Suppose not, and let $G$ be a (bull, $P_{6},F$)-free MNWD graph. By Lemma 2.3, we have that $G$ is a connected odd net-free graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect and contains no odd antihole of number of vertices at least 7. By Lemmas 2.5 and 2.6, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Theorem 1.1, we have that $G[M(v)]$ contains a 5-hole. Therefore, by Lemma 4.1, $G$ has a homogeneous set, which contradicts with Lemma 2.1. This completes the proof of Theorem 1.3. $\square$
5 Proof of Theorems 1.5, 1.6 and 1.7
In this section, we will prove Theorems 1.5-1.7. Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$. Let $\cal C$} be a class of graphs. We denote by ${\cal C}_{s}$ the closure of $\cal C$} under taking substitution.
Lemma 5.1
[…]
Lemma 5.2
[…]
Lemma 5.3
((Theorem 6.1 of [2] )) Let $\cal C$} be a class of graphs. If $\cal C$} is hereditary and polynomially $\chi$-bounded with $\chi\leq\omega^{k}$, then ${\cal C}_{s}$ is polynomially $\chi$-bounded with $\chi\leq\omega^{2k+3}$.
Lemma 5.4
((Lemma 5 of [4])) Let $G$ be a connected (bull, odd net)-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
Lemma 5.5
[17] The class of perfect graphs is closed under taking substitution.
We now proceed to prove Theorems 1.5-1.7.
Proof of Theorem 1.5: We prove the Theorem by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. If $\omega(G)\leq 2$, by Lemma 2.5, then $\chi(G)\leq 4$ and we are done. Therefore, by inductive hypothesis, we may assume that
$G$ is connected and $\omega(G)\geq 3$. (6)
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ is perfect, then $(N(v),M(v)\cup\{v\})$ is a partition of $V(G)$ which satisfies the lemma since $G$ is odd net-free. Therefore, we may assume that $G[M(v)]$ is imperfect. By Lemma 5.4, $G[M(v)]$ contains no odd antihole of number of vertices at least 7. So, by Theorem 1.1, $G[M(v)]$ contains a 5-hole as $G$ is $P_{6}$-free. By Lemma 4.1, we have that $G$ has a homogeneous set, say $S$. Recall that $1<|S|<|V(G)|$.
Since $G$ is connected, we have that there exists a vertex which is complete to $S$, and so $G[S]$ is perfect as $G$ is odd net-free. Since $S\neq\mbox{{\rm\O }}$, we may assume that $x\in S$. Let $G^{\prime}=G[V(G)\setminus(S\setminus\{x\})]$. Therefore, by inductive hypothesis, there exists a partition $(X^{\prime},Y^{\prime})$ of $V(G^{\prime})$ such that $G[X^{\prime}]$ and $G[Y^{\prime}]$ is perfect. Since $x\in V(G^{\prime})$, without loss of generality, we may assume that $x\in X^{\prime}$. Let $X=X^{\prime}\cup S$ and $Y=Y^{\prime}$. Clearly, $(X,Y)$ is a partition of $V(G)$. Since $S$ is a homogeneous set of $G$, we have that $S$ is a homogeneous set of $G[X]$, and thus $G[X]$ is a graph obtained from $G[X^{\prime}]$ by substituting the vertex $x$ by the graph $G[S]$. By Lemma 5.5, $G[X]$ is perfect. Therefore, the partition $(X,Y)$ satisfies the lemma. This completes the proof of Theorem 1.5. $\square$
Proof of Theorem 1.6: We prove the Theorem by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. We may assume that $\omega(G)\geq 2$ and $G$ is connected by inductive hypothesis.
If $G$ is odd net-free, then $\chi(G)\leq 2\omega(G)\leq\omega(G)^{2}$ by Theorem 1.5, and so we are done. Therefore, we may assume that $G$ contains an odd net. Now, by Lemma 5.2, there exists a vertex $v$ such that $G[M(v)]$ is perfect. By inductive hypothesis, we have that $\chi(G[N(v)])\leq(\omega(G)-1)^{2}$. Therefore, we have that
$\chi(G)\leq\chi(G[N(v)])+\chi(G[\{v\}\cup M(v)])\leq(\omega(G)-1)^{2}+\omega(G)\leq\omega(G)^{2}.$
This completes the proof of Theorem 1.6. $\square$
Proof of Theorem 1.7: Let $\cal G$} be the class of (bull, $P_{6}$)-free graphs, and $\cal C$} be the class of (bull, $P_{6}$, odd net+)-free graphs. By Lemma 5.3 and Theorem 1.6, it suffices to show that ${\cal G}\subseteq{\cal C}_{s}$.
Let $G\in{\cal G}$. Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a (bull, odd net+)-free graph $G_{1}$ by substituting a vertex by (bull, odd net+)-free -free graph $G_{2}$. Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraphs of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free, which implies that $G_{1},G_{2}\in{\cal C}$. Therefore, $G\in{\cal C}_{s}$, and so ${\cal G}\subseteq{\cal C}_{s}$.
This completes the proof of Theorem 1.7. $\square$
Declarations
  • Ran Chen is supported by Postgraduate Research and Practice Innovation Program of Jiangsu Province KYCX25_1926. Di Wu is supported by the Scientific Research Foundation of Nanjing Institute of Technology, China (No. YKJ202448).
  • Conflict of interest The authors declare no conflict of interest.
  • Data availibility statement This manuscript has no associated data.